If a homeowner places the top of a twenty foot ladder against the house and twelve feet above the ground, how far from the home is the base of the ladder?
I do not know how to set up the problem, please help.

Answer by rfer(16322)

a^2+b^2=c^2
----------
b^2=c^2-a^2
b^2=20^2-12^2
b^2=400-144
b^2=256
sqrt 256=16 ft